A mixture of He, Ne, and Ar has a pressure of 12.4 atm at 28.0 °C. If the partial pressure of He is 4.49 atm and that of Ar is 2.77 atm, what is the partial pressure of Ne?
Answer #1

Answer

Ptotal= Phelium+Pneon+Pargon

Pneon=Ptotal-(Phelium+Pargon)

Pneon= 12.4-(4.49+2.77)

Pneon= 12.4-(7.26)= 5.14 atm